# On the Degrees of Freedom of some Lasso procedures

### v1

@@ Appendix A Proofs
If the increment of $\mathbf{y}$ is small enough, e.g. $|\Delta\mathbf{y}|<\varepsilon$, zou2007 showed that for Lasso the projection matrix $\mathbf{H}_{\gamma}$ and the function $\eta_{\gamma}$ remain constant. Indeed, if $\gamma$ is not a transition point, small perturbations of $\mathbf{y}$ does not affect neither $\mathbf{H}_{\gamma}$ nor $\eta_{\gamma}$. In our case, instead, we have that $\mathbf{H}_{\gamma}(\mathbf{y}+\Delta\mathbf{y})=\mathbf{H}_{\gamma}(\mathbf{y})$ but $\eta_{\gamma}(\mathbf{y}+\Delta\mathbf{y})\neq\eta_{\gamma}(\mathbf{y})$, because $\mathbf{W}_{\mathcal{A}}(\mathbf{y}+\Delta\mathbf{y})\neq\mathbf{W}_{\mathcal{A}}(\mathbf{y})$. Therefore
@@
which is the result presented in yuan:2006:glasso. By contrast,
@@
which is the result presented in vaiter_etal.2017. Similar simplifications under the non orthonormal setting are not easy to derive, as this proof is based on the eigenvalues of the matrix $\mathbf{B}$ which are unknown unless $\mathbf{X}^{\top}\mathbf{X}=\mathbf{I}_{p}$. ∎
@@
By exploiting the properties of the vectorization operator, as detailed in magnus_neudecker.1999, we get:

### v2

@@ Appendix A Proofs
If the increment of $\mathbf{y}$ is small enough, e.g. $|\Delta\mathbf{y}|<\varepsilon$, Zou et al. (2007) showed that for Lasso the projection matrix $\mathbf{H}_{\gamma}$ and the function $\eta_{\gamma}$ remain constant. Indeed, if $\gamma$ is not a transition point, small perturbations of $\mathbf{y}$ does not affect neither $\mathbf{H}_{\gamma}$ nor $\eta_{\gamma}$. In our case, instead, we have that $\mathbf{H}_{\gamma}(\mathbf{y}+\Delta\mathbf{y})=\mathbf{H}_{\gamma}(\mathbf{y})$ but $\eta_{\gamma}(\mathbf{y}+\Delta\mathbf{y})\neq\eta_{\gamma}(\mathbf{y})$, because $\mathbf{W}_{\mathcal{A}}(\mathbf{y}+\Delta\mathbf{y})\neq\mathbf{W}_{\mathcal{A}}(\mathbf{y})$. Therefore
@@
which is the result presented in Yuan and Lin (2006). By contrast,
@@
which is the result presented in Vaiter et al. (2017). Similar simplifications under the non orthonormal setting are not easy to derive, as this proof is based on the eigenvalues of the matrix $\mathbf{B}$ which are unknown unless $\mathbf{X}^{\top}\mathbf{X}=\mathbf{I}_{p}$. ∎
@@
By exploiting the properties of the vectorization operator, as detailed in Magnus and Neudecker (1999), we get: